Suppose the following demand and supply function:

Qd = 750 – 25P

Qs = -300 + 20 P


       i.           Find equilibrium price and quantity

     ii.           Find consumer and producer surplus



1
Expert's answer
2021-04-29T18:56:19-0400

(I)

At equilibrium Qd=Qs

"750 \u2013 25p=-300+20p"

"750+300=20p+25p"

"1050=45p"

"P=23.33"

Substituting value of P in demand function

"=750-25(23.33)"

"=750-583.25"

"Q=166.7"

(ii)

Graphically the consumer and producer can be represented as



The consumer and producer surplus can be calculated with the formula:

Consumers|producers surplus=1/2× Base × Height


According to the graph mentioned above, Ps and PD is also calculated. At both these price points quantity demanded and supplied are 0. 

"0=750-25P_D"

"25P_D=750"

"P_D=\frac{750}{25}"

"=30"


"0=-300+25P_S"

"20P_S=300"

"P_S=\frac{300}{20}"

"=15"

consumer and producer surplus are then calculated as

Consumer surplus"=\frac{1}{2}×166.7×(30-23.33)"

"=\frac{1}{2}×166.7×6.67"

"=555.94"

Producer surplus

"=\frac{1}{2}×166.7×(23.33-15)"

"=\frac{1}{2}×166.7×8.33"

"=694.31"